OpenStudy (anonymous):

The sum of the ages of a woman and her son is 36 years. Six years from now the woman will be twice as old as her son. How old is she now?

OpenStudy (ahsome):

Use algebra to answer Let S be Son's Age Let W be Woman's age S+W=36 W+6=2S

OpenStudy (ahsome):

You can then to substitution or subtraction

OpenStudy (anonymous):

i dont understand. it has 2 equations?

OpenStudy (ahsome):

If you see, it says Son and Woman is 36 years. Therefore S + W = 36, right? Then it says that when the woman's age is plus 6, then it would be twice the son. Therefore Woman + 6 = 2 X Son, W+6=2S

OpenStudy (anonymous):

okay. thank you

OpenStudy (ahsome):

No problem :D

OpenStudy (anonymous):

i have to solve it too...

OpenStudy (ahsome):

Whoops, let me try again Equation is: \[s+w=36\] & \[w+6=2s\] Rearrange so it is \[w+s=16\] & \[w+6=2s\] Subtract each other. The W would cancel out \[s-6=16-2s\]

OpenStudy (anonymous):

okay i get it now

OpenStudy (ahsome):

:D

OpenStudy (anonymous):

where did the 16 come from?

OpenStudy (ahsome):

Oh, whoops 36. Computer is not working ;)

OpenStudy (anonymous):

so s=14 and w=22 but that equals 36 so now i need to do it for six years from now

OpenStudy (ahsome):

22+6

OpenStudy (ahsome):

Thats how old she would be in 6 years

OpenStudy (anonymous):

Thank you! I did it! :)

OpenStudy (ahsome):

If you want to make sure, type "s+2=36, w+6=ws" into WolframAlpha. That will give the answers for you to chec

OpenStudy (ahsome):

No Problem :D